OpenStudy (anonymous):

Solve the proportion: 4/(x+1) = 3/(x+2)

OpenStudy (anonymous):

how would i do this please explain

OpenStudy (anonymous):

\[\frac{ 4 }{ (x+1) }=\frac{ 3 }{ (x+2) }\] do cross multiplication

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

its -5

OpenStudy (anonymous):

right

OpenStudy (whpalmer4):

Try substituting your answer into the original equation. Do you get a true statement if you do that? If you do not, then your answer cannot be correct.